THE FOURIER TRANSFORM MICROWAVE SPECTRUM AND STRUCTURE OF OPF

BETHANY GATEHOUSE, Th. BRUPBACHER and \underline{MICHAEL C. L. GERRY}, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver B.C., Canada V6T 1Z1.

The pure rotational spectrum of free gas phase OPF has been observed in the frequency range 4-26 GHz, using a cavity pulsed jet Fourier transform microwave spectrometer. The samples were prepared in mixtures of PF_{3} and O_{2} in Ne, using a pulsed discharge nozzle. The transitions were analysed in terms of rotational, centrifugal distortion and nuclear spin-rotation coupling constants. The determined rotational constants have been used to evaluate r_{0}, r_{s} and approximate r_{e} geometries, which are compared to those of related molecules. The spin-rotation coupling constants have been used to calculate the nuclear shielding parameters of the P and F nuclei; however, because both nuclei have I = \frac{1}{2} , and because both sets of coupling constants have very similar values, assignment to a given nucleus was not possible, based on experimental data alone.